What is the probability of getting heads when flipping a coin and getting a number?
Originally Answered: what is the probability of getting heads when flipping a coin and getting a number greater than or equal to 5 when rolling a single die? The probability of getting heads is 1/2, and the probability of getting 5 or 6 is 1/3, and so you simply multiply 1/2 x 1/3, which is 1/6.
What is the probability that you flip a coin landing on heads and roll a dice landing on an odd number?
Probability for tossing on heads=0.5 Probability of rolling on odd number on die (1 or 3 or 5)=0.5 As per addition rule (A union B, A or B) that is 0.5+0.5=1 that seems impossible. How could it be? You could very well get a tail and an even-numbered die.
What is the probability of flipping a coin and it landing on heads 5 times in a row?
Your proposed answer of 13/32 is correct. If there are four or five heads in the sequence of five coin tosses, at least two heads must be consecutive.
What is the probability of getting 2 heads and 2 tails?
1/4
For example, the probability of two heads is 1/2 · 1/2 = 1/4, and the probability of two tails is the same. Since there are two leaves corresponding to one head and one tail, each of probability 1/4, the probability of this event is 1/4 + 1/4 = 1/2.
What is the probability of getting one head and two tails?
Summary: The Probability of getting two heads and one tails in the toss of three coins simultaneously is 3/8 or 0.375.
How do you calculate the probability of flipping a coin?
What Are Coin Toss Probability Formulas?
- On tossing a coin, the probability of getting head is: P(Head) = P(H) = 1/2.
- Similarly, on tossing a coin, the probability of getting a tail is: P(Tail) = P(T) = 1/2.
What is the probability of 5 coins?
The probability of 5 heads out of 5 coins is (1/2)^5 or 1/32.
How do you solve a probability coin problem?
When we toss a coin the possible outcomes are?
two
The two outcomes of the toss of a coin are heads or tails. For any individual toss of the coin, the outcome will be either heads or tails. The two outcomes (heads or tails) are therefore mututally exclusive; if the coin comes up heads on a single toss, it cannot come up tails on the same toss.
What is the probability of tossing 3 coins?
⅛
Solution: When 3 coins are tossed, the possible outcomes are HHH, TTT, HTT, THT, TTH, THH, HTH, HHT. (i) Let E1 denotes the event of getting all tails. Hence the required probability is ⅛.